OpenStudy (anonymous):

Write the equation of the line that passes through the points (–1, –4) and (2, 5).

OpenStudy (anonymous):

First of all, you should find the slope. Do you know how to do that?

OpenStudy (anonymous):

no, i dont

OpenStudy (anonymous):

Do you know the formula?

OpenStudy (anonymous):

y=mx+b?

OpenStudy (anonymous):

That's the slope intercept form. To find the slope, you would need the formula Slope=(y-y1)/(x-x1)

OpenStudy (anonymous):

ohh allright

OpenStudy (anonymous):

Find the slope for me, will you?

OpenStudy (anonymous):

i will

OpenStudy (anonymous):

People get annoyed when I give out answers AND work so I try to make you do work too -_-;;

OpenStudy (anonymous):

lol its ok, the slope is -6

OpenStudy (anonymous):

errr. I think the slope is 3. (5-(-4))/(2-(-1))?

OpenStudy (anonymous):

Anyway, you know that the slope intercept form is y=mx+b with the x and y as coordinates, the b being a random constant, and the m being the slope. So you know that the slope is 3 now, so you can plug it in the equation: y=3x+b You can then plug in on of your coordinates sets ( cuz you have two). I like to stay with positive numbers so I would do this: 5=3(2)+b You do that equation to get b. Tell me what you get.

OpenStudy (anonymous):

b=-1?

OpenStudy (anonymous):

Yesss. Now what you do is keep your slope and b constant annd throw away the x and y because the x and y aren't constant, or else the equation would be of one point. So you take y=mx+b and plug in m and b to get the equation of the line

OpenStudy (anonymous):

thank you soooo much!<3

OpenStudy (anonymous):

Just in case, what's your final answer?